Edwin conducted a survey to find the percentage of people in an area who smoke regularly. He defined the label smoking regularly for males smoking 30 more cigarets in A-day and for female smoking 20 or more. Out of 635 persons who took part in the survey 71 are labeled as people who smoke regularly. What is the standard error for the sample proportion?
Answer #1

Sample size (n) =635

Favourable case = 71

Sample proportion (p) = 71/635 =

Standard error = ✓[p*(1-p))/n] = ✓[ (71/635 * 564/635)/635] = 0.012505703

Dear if you are satisfied with my explanations please THUMBS UP... Thanks in advance for your support and love...